# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

### v1

@@ Abstract
Using bidifferential calculus, we derive a vectorial binary Darboux transformation for an integrable matrix version of the first negative flow of the Kaup-Newell hierarchy. A reduction from the latter system to an integrable matrix version of the Fokas-Lenells equation is then shown to inherit a corresponding vectorial Darboux transformation. Matrix soliton solutions are derived from the trivial seed solution. Furthermore, the Darboux transformation is exploited to determine in a systematic way exact solutions of the two-component vector Fokas-Lenells equation on a plane wave background. This includes breathers, dark solitons, rogue waves and “beating solitons”.
@@ 1 Introduction
where $u$ is a complex function of independent real variables $x$ and $t$, and a subscript $x$ or $t$ indicates a partial derivative with respect to $x$, respectively $t$. In particular, it models propagation of ultrashort light pulses in birefringend optical fibers, taking certain nonlinear effects into account. For the integrability features of (1.1), see [4] and the references cited there. The FL equation (1.1) possesses a generalization to integrable matrix versions
@@
Here $u$ is now a matrix of any size, where the components are functions of $x$ and $t$, $u^{\dagger}$ denotes the conjugate transpose of $u$, and $\sigma_{1},\sigma_{2}$ are any constant involutary Hermitian matrices of suitable size. In the case, where $u$ is a vector, (1.2) already appeared in [5] (see (82) therein), also see [6, 7, 8, 9, 10, 11, 12, 13]for the two-component vector case, and [14, 15] for the three-component case. Some results on the $2\times 2$ matrix case have been obtained in [15].
@@
In this work, we present a vectorial Darboux transformation for (1.2), which allows to generate exact solutions from a given “seed” solution $u$. It involves a linear system that depends on $u$ in such a way that its compatibility (or integrability) condition is satisfied if $u$ solves (1.2), and the dependent variables of the linear system are $n$-component vectors, where $n$ corresponds to the number of solitons in the generated class of solutions of (1.2). For $n=1$, the linear system reduces to a Lax pair, depending on a “spectral parameter”.
@@
Though essentially self-contained, this work is based on [4] (which, in turn, partly draws from [16]), where a vectorial Darboux transformation for the scalar FL equation has been derived, using bidifferential calculus [17] (also see, in particular, [18, 19]). In Section 2 we show that, by choosing a suitable (realization of the) bidifferential calculus, a simple “Miura transformation” equation (see (2.1)) is essentially turned into a matrix generalization of the first “negative flow” of the Kaup-Newell (KN) hierarchy,
@@
The crucial point is that the Miura transformation equation possesses a (vectorial) binary Darboux transformation (bDT). A bDT allows to generate from a given solution a new family of solutions, using a linear system and also another (“adjoint”) linear system. The bDT for (1.3) is derived in Section 3.
A difficult task is to find an extension of the reduction (1.4) such that the bDT generates solutions of the matrix FL equation (1.2), i.e., that it preserves the reduction condition (1.4). This problem is solved in Section 4, essentially following the steps in the scalar case [4]. In Section 4 we also derive (multi-) soliton solutions from the trivial solution $u=0$.
Section 5 deals with the vector version of the matrix FL equation, obtained when the matrix $u$ in (1.2) reduces to a vector (also see [5, 6, 7, 8, 9, 14, 10, 12, 13]). In Section 6, we finally concentrate on the two-component vector case. The vectorial Darboux transformation is used to generate various types of soliton solutions on a plane wave background solution.
Whereas, in the classical (single step) Darboux transformation, dark solitons cannot be obtained directly, but only indirectly via a certain limit procedure (see [7, 8] for the case of the two-component vector FL equation), the vectorial Darboux transformation offers a direct way. A key feature of a vectorial Darboux transformation is the appearance of a Sylvester equation for the so-called Darboux potential $\Omega$. When it possesses a unique solution, the generated solution of the respective nonlinear integrable equation can also be reached by iteration of a classical Darboux transformation (using different spectral parameters). But the important class of dark solitons of NLS-type equations is obtained via the complementary case, where the Sylvester equation does not (completely) determine the Darboux potential and imposes constraints on its right hand side, and thus on the solution of the respective linear system. Vectorial Darboux transformations obtained via bidifferential calculus come along with additional equations that directly determine the Darboux potential in the latter case. This is a systematic procedure in contrast to the limit technique of what is sometimes called “generalized Darboux transformation”, which involves a kind of fine-tuning of parameters.
The “non-recursive Darboux transformation” in [20], applied in [10] to the two-component vector FL equation, is also “vectorial”, but lacks the important step of promoting the spectral parameter to a matrix and thus misses the superposition rule for solutions with the same seed (cf. Remark 4.4 below).
Finally, Section 7 contains some concluding remarks. By a solution of a differential equation, throughout this work we mean a strong solution, with differentiability class required by the equation. For example, $u$ in (1.1) - (1.3) has to be of class $\mathcal{C}^{2}$.
@@
In the following, we show that the matrix versions (1.3) of the first negative flow of KN hierarchy arise from the simple equation
@@
Let $(u,v)$ be a solution of (1.3). Then
@@
The linear first order matrix ODEs for $g_{22}$ and $S$ possess invertible solutions (see [22], for example). The statement of the lemma is then easily verified. ∎
@@
Elaborating this result for the choice of bidifferential calculus in Section 2, leads to the following theorem.
@@
with any constant invertible $n\times n$ matrices $S$ and $T$. As a consequence, without restriction of generality, we may assume that $\Delta$ and $\Gamma$ are in (possibly different, upper- or lower-triangular) Jordan normal form. $\Box$
@@ 4 A vectorial Darboux transformation for the matrix FL equation
Let $\Gamma$ be an invertible constant $n\times n$ matrix, such that $\Gamma$ and $\Gamma^{\dagger}$ have no common eigenvalue. Let $u$ be a solution of the $m_{2}\times m_{1}$ matrix Fokas-Lenells equation (1.2) on an open set $\mathcal{U}\subset\mathbb{R}^{2}$. Furthermore, let $\eta_{1}$ and $\tilde{\eta}_{2}$ be $n\times m_{1}$, respectively $n\times m_{2}$, matrix solutions of the linear system
@@
If $\Gamma^{(1)}$ and $\Gamma^{(2)\dagger}$ have no common eigenvalue, these equations have unique solutions $\Omega^{(12)}$ and $\Omega^{(21)}$. Otherwise additional equations, obtained from (4.5) and (4.6), have to be solved. $\Box$
@@ 4.1 Zero seed solutions
For $u=0$ (zero seed), (4.1) is solved by
@@
Assuming the spectrum condition for $\Gamma$, the Lyapunov equation possesses a unique solution. According to Theorem 4.1,
@@
solves the matrix FL equation (1.2).
@@
$\Omega_{ij}=\frac{1}{\gamma_{i}+\gamma_{j}^{\ast}}\,\Big{(}\mathrm{i}\,\boldsymbol{a}_{ij}\,\gamma_{j}^{\ast}\,e^{-\varphi(\gamma_{i})-\varphi(\gamma_{j}^{\ast})}+\boldsymbol{b}_{ij}\,e^{\varphi(\gamma_{i})+\varphi(\gamma_{j}^{\ast})}\Big{)}\,$,
where
$\varphi=\frac{1}{2}(\gamma^{-1}\,x+\gamma\,t)\,.$ (4.9)
@@
$k_{i}\,\boldsymbol{a}_{ii}-\boldsymbol{b}_{ii}=0\,,\qquad i=1,\ldots,n\,.$
@@
$u^{\prime}=-\frac{k\,e^{-\mathrm{i}\,k^{-1}(x-k^{2}t)}}{\boldsymbol{a}\,(x+k^{2}t)+\mathrm{Re}(c_{1})\,k+\frac{1}{2}\mathrm{i}\,\boldsymbol{a}\,k}\,\sigma_{2}\,a_{2}^{\dagger}\,a_{1}\,$,
@@
where $\varphi$ is again given by (4.9), and
$\rho=\frac{1}{2}\big{(}\gamma^{-2}x-t\big{)}\,.$
After having determined
@@
If $\kappa=2\mathrm{Re}(\gamma)\neq 0$, the solution of the Lyapunov equation is given by
@@
Besides the exponential dependence via $A_{ij}$ and $B_{ij}$, the solution (4.15) also depends on the independent variables $x$ and $t$ through the linear expression $\rho$. Generically, the determinant of $\Omega$ has non-zero real and imaginary parts, so that $u^{\prime}$ is regular. We obtain matrix generalizations of “positon” solutions, see [4] for the scalar case.
If $\gamma$ is imaginary, i.e., $\gamma=-\mathrm{i}\,\,k$ with $k\in\mathbb{R}\setminus\{0\}$, then we have $\rho=-\frac{1}{2}(x/k^{2}+t)=\rho^{\ast}$ and the Lyapunov equation requires
@@
$\mathrm{Im}(a_{21}\,a_{22}^{\ast})-\sigma_{1}\sigma_{2}\,\Big{(}\frac{1}{2}|a_{11}|^{2}-k\,\mathrm{Im}(a_{11}\,a_{12}^{\ast}\Big{)}=0\,.$ (4.19)
@@
so that $\mathrm{Re}(c_{22})$ and $\mathrm{Im}(c_{12})$ remain as free parameters. Let $k$ be determined by (4.18). The constants $a_{ij}$, $i,j=1,2$, are still constrained by (4.19). Then (4.15) is a quasi-rational solution of the $m_{2}\times m_{1}$ matrix FL equation. If $\boldsymbol{a}_{11}\neq 0$, for constant $t$ (respectively $x$) the numerators of components of $u^{\prime}$ grow at most with $x^{3}$ (respectively $t^{3}$), whereas $\det(\Omega)$ grows with the fourth power. On a line, along which $\rho$ and thus $x+k^{2}t$ is constant, $|u^{\prime}|\sim 1/|t|$ as $t\to\pm\infty$. As a consequence, the solution must be localized along a worldline, which is not a straight line, a property attributed to “positons”. In particular, our results supplement the treatment of the scalar FL equation ($m_{1}=m_{2}=1$) in Section 4.1 of [4], also see Fig. 1 for two examples.

### v2

@@ Abstract
Using bidifferential calculus, we derive a vectorial binary Darboux transformation for an integrable matrix version of the first negative flow of the Kaup-Newell hierarchy. A reduction from the latter system to an integrable matrix version of the Fokas-Lenells equation is then shown to inherit a corresponding vectorial Darboux transformation. Matrix soliton solutions are derived from the trivial seed solution. Furthermore, the Darboux transformation is exploited to determine in a systematic way exact solutions of the two-component vector Fokas-Lenells equation on a plane wave background. This comprises breathers, dark solitons, rogue waves and “beating solitons”.
@@ 1 Introduction
where $u$ is a complex function of independent real variables $x$ and $t$, and a subscript $x$ or $t$ indicates a partial derivative with respect to $x$, respectively $t$. In particular, it models propagation of ultrashort light pulses in birefringend optical fibers, taking certain nonlinear effects into account. For some integrability features of (1.1), see [4] and the references cited there. The FL equation (1.1) possesses a generalization to integrable matrix versions
@@
Here $u$ is now a matrix of any size, where the components are functions of $x$ and $t$, $u^{\dagger}$ denotes the conjugate transpose of $u$, and $\sigma_{1},\sigma_{2}$ are any constant involutary Hermitian matrices of suitable size. In the case, where $u$ is a vector, (1.2) already appeared in [5] (see (82) therein), also see [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] for the two-component vector case, and [17] for the three-component case. Some results on the $2\times 2$ matrix case have been obtained in [18].
@@
In this work, we present a vectorial Darboux transformation for (1.2), which allows to generate exact solutions from a given “seed” solution $u$. It involves a linear system that depends on $u$ in such a way that its compatibility (or integrability) condition is satisfied if $u$ solves (1.2), and the dependent variables of the linear system are $n$-component vectors, where $n$ may be thought of as corresponding to the number of solitons in the generated class of solutions of (1.2) (though already some $n=1$ solutions can be regarded as being composed of more elementary solutions). For $n=1$, the linear system reduces to a Lax pair, depending on a “spectral parameter”. In our linear system, this parameter is promoted to a “spectral matrix”.
The original Darboux transformation [19] has to be iterated in order to generate multiple soliton solutions of an integrable equation. Moreover, in each step a different spectral parameter value has to be used. Accordingly, important classes of solutions are only obtained indirectly by taking special limits, where a priori different spectral parameters tend to the same value. This concerns “multiple pole solutions” (in the language of the inverse scattering method), which include “positons” of the Korteweg-deVries and related equations, for which V.B. Matveev set up a generalized Darboux transformation [20, 21]. Later a corresponding generalization of the classical Darboux transformation has also been formulated for the Nonlinear Schrödinger (NLS) and related integrable equations in order to generate higher order rogue waves and (multi-) dark soliton solutions (see, e.g., [22, 23], and [7, 8] for applications to the case of the two-component vector FL equation).
A vectorial Darboux transformation does not require iteration in order to generate multi-soliton solutions, but does the job in a single step. Furthermore, it offers a direct way to the abovementioned important classes of solutions. [^1] The “Cauchy matrix approach”, applied in [24] to the scalar FL equation, may be regarded as a vectorial Darboux transformation restricted to the case of trivial seed. Moreover, it allows us, at least to some extent and in a tractable way, to exhaust the various possibilities of exact solutions within this framework. A beautiful aspect of a vectorial Darboux transformation is a simple superposition rule on the level of data (consisting of a “spectral matrix” and the corresponding solution of the linear system) [^2] In case of a vectorial binary Darboux transformation, we are dealing with two spectral matrices and two linear systems. determining a solution of the integrable nonlinear equation for a given seed solution, see Remark 4.4 below.
The usefulness of a vectorial version of the classical Darboux transformation has also been demonstrated by the “non-recursive Darboux transformation” in [25], which has been applied in [11] to construct higher order rogue wave solutions of the two-component vector FL equation. It should be noted, however, that it lacks the, in our opinion important, step of promoting the spectral parameter more consequently to a matrix and thus misses the superposition rule for solutions with the same seed. In any case, the “non-recursive Darboux transformation” is a step towards a more complete vectorial Darboux transformation, as proposed earlier by Mañas [26] for the NLS equation.
A key feature of a vectorial Darboux transformation is the appearance of a Sylvester equation for the so-called Darboux potential $\Omega$. The input for this algebraic matrix equation is the spectral matrix and a matrix built from the solution of the linear system. When the spectral matrix is diagonal and the Sylvester equation has a unique $n\times n$ matrix solution, it is expected that the generated solution of the respective nonlinear integrable equation can also be reached by an $n$-fold iteration of a classical Darboux transformation (using different spectral parameters). Multiple pole solutions and higher order rogue waves are obtained by choosing the spectral matrix as a Jordan block, and the nature of the generated class of solutions depends on (properties of) its eigenvalue.
Important classes of solitons of NLS-type equations are obtained, however, via the complementary case, where the Sylvester equation does not (completely) determine the Darboux potential and imposes constraints on the solution of the respective linear system. [^3] That the solution of the linear system can indeed be reduced in such a way that these constraints are satisfied, appears to be a bit of a miracle. Vectorial Darboux transformations, like those obtained via bidifferential calculus [27] (also see, in particular, [28, 29]), come along with additional differential equations that completely determine the Darboux potential in such a “degenerate case”. Corresponding solution classes will be worked out explicitly for the two-component vector FL equation.
In the present work, we exploit the vectorial Darboux transformation (only) with seed solutions, for which the linear system can be turned into a system of linear PDEs with constant coefficients. Then, generically, the general solution is a linear combination of exponentials of a linear expression in $x$ and $t$. But when the corresponding characteristic equation has a multiple root, the linear system “degenerates” and the general solution also involves a polynomial (or polynomials) of the independent variables. If the seed solution is an exponential describing a plane wave, and if we have a root of maximal order, then the resulting solution of the vector FL equation (or, more generally, some NLS-type equation) is quasi-rational, i.e., a rational expression multiplied by the plane wave exponential. If the norm of such a solution is localized on the uniform background, it represents a rogue wave (or freak wave), a case that attracted enormous interest, since it describes a wave that suddenly appears (“from nowhere”) and then disappears again.
There is another, independent mechanism, how rational expressions can enter the solution stage. This is at work in those cases, where the Sylvester equation does not possess a unique solution, which is true if and only if the spectral matrix does not satisfy the so-called spectrum condition.
@@
Though essentially self-contained, this work is based on [4] (which, in turn, partly draws from [24]), where a vectorial Darboux transformation for the scalar FL equation has been derived, using bidifferential calculus. In Section 2 we show that, by choosing a suitable (realization of the) bidifferential calculus, a simple “Miura transformation” equation (see (2.1)) is essentially turned into a matrix generalization of the first “negative flow” of the Kaup-Newell (KN) hierarchy,
@@
The crucial point is that the Miura transformation equation possesses a vectorial binary Darboux transformation (bDT). A bDT allows to generate from a given (“seed”) solution a new family of solutions, using solutions of a linear system and also of another (“adjoint”) linear system. The bDT for (1.3) is derived in Section 3.
A difficult task is to find an extension of the reduction (1.4) such that the bDT generates solutions of the matrix FL equation (1.2), i.e., that it preserves the reduction condition (1.4). This problem is solved in Section 4, essentially following the steps in the scalar case [4]. In Section 4 we also derive (multi-) matrix soliton solutions from the trivial solution $u=0$.
Section 5 deals with the vector version of the matrix FL equation, obtained when the matrix $u$ in (1.2) reduces to a vector (also see [5, 6, 7, 8, 9, 17, 11, 14, 16]). In Section 6, we finally concentrate on the two-component vector case. The vectorial Darboux transformation is used to generate all types of soliton solutions on a plane wave background solution in a systematic way.
Finally, Section 7 contains some concluding remarks.
@@
In the following, we show that the matrix versions (1.3) of the first negative flow of the KN hierarchy arise from the simple equation
@@
Let $(u,v)$ be a solution [^4] A solution may not be defined everywhere on $\mathbb{R}^{2}$. of (1.3). Then
@@
The linear first order matrix ODEs for $g_{22}$ and $S$ possess invertible solutions (see [31], for example), wherever $u$ and $v$ are smooth. The statement of the lemma is then easily verified. ∎
@@
Elaborating this result for the choice of bidifferential calculus in Section 2, leads to the following theorem. [^5] Here we restrict our considerations to a framework of finite-dimensional matrices. But under suitable technical assumptions, we may let $n$ in Theorem 3.2 to be infinite, or better generalize the matrices involving the matrix size $n$ to become (bounded) operators on a Banach space. Also see [26], for example.
@@
with any constant invertible $n\times n$ matrices $S$ and $T$. As a consequence, without restriction of generality, we may assume that $\Delta$ and $\Gamma$ are in (possibly different, upper or lower triangular) Jordan normal form. $\Box$
@@ 4 A vectorial Darboux transformation for the matrix FL equation
Let $\Gamma$ be an invertible constant $n\times n$ matrix, such that $\Gamma$ satifies the “spectrum condition”, i.e., $\Gamma$ and $\Gamma^{\dagger}$ have no common eigenvalue. Let $u$ be a solution of the $m_{2}\times m_{1}$ matrix Fokas-Lenells equation (1.2) on an open set $\mathcal{U}\subset\mathbb{R}^{2}$. Furthermore, let $\eta_{1}$ and $\tilde{\eta}_{2}$ be $n\times m_{1}$, respectively $n\times m_{2}$, matrix solutions of the linear system
@@
If $\Gamma^{(1)}$ and $\Gamma^{(2)\dagger}$ have no common eigenvalue, these equations possess unique solutions $\Omega^{(12)}$ and $\Omega^{(21)}$. Otherwise additional equations, obtained from (4.5) and (4.6), have to be solved. $\Box$
@@ 4.1 Zero seed solutions
For $u=0$ (zero seed), (4.1) is completely solved by
@@
If the spectrum condition for $\Gamma$ holds, the Lyapunov equation possesses a unique solution $\Omega$. If not, then the Lyapunov equation leads to constraints and we also have to use the equations in Corollary 4.2 to determine $\Omega$. According to Theorem 4.1 and Corollary 4.2,
@@
solves the matrix FL equation (1.2). In the following, we set
$\varphi(\Gamma)=\frac{1}{2}(\Gamma^{-1}\,x+\Gamma\,t)\,$,
and work out some classes of solutions.
@@
$\Omega_{ij}=\frac{1}{\gamma_{i}+\gamma_{j}^{\ast}}\,\Big(\mathrm{i}\,\boldsymbol{a}_{ij}\,\gamma_{j}^{\ast}\,e^{-\varphi(\gamma_{i})-\varphi(\gamma_{j}^{\ast})}+\boldsymbol{b}_{ij}\,e^{\varphi(\gamma_{i})+\varphi(\gamma_{j}^{\ast})}\Big)\,.$
@@
$\boldsymbol{b}_{ii}=k_{i}\,\boldsymbol{a}_{ii}\,,\qquad i=1,\ldots,n\,.$
@@
$u^{\prime}=-\frac{e^{-\mathrm{i}\,\big(k^{-1}x-k\,t\big)}}{\boldsymbol{a}\,(k^{-1}x+k\,t)+\mathrm{Re}(c_{1})+\frac{1}{2}\mathrm{i}\,\boldsymbol{a}}\,\sigma_{2}\,a_{2}^{\dagger}\,a_{1}\,$,
@@
Example 4.7.
Let $\Gamma=-\mathrm{i}\,k\,I_{n}$, where $k\in\mathbb{R}\setminus\{0\}$. Then the Lyapunov equation becomes the constraint $\boldsymbol{b}=k\,\boldsymbol{a}$, and (4.6) leads to $\Omega=(k^{-1}x+k\,t)\,\boldsymbol{a}+\boldsymbol{c}$, with a constant $n\times n$ matrix $\boldsymbol{c}$, for which (4.5) requires $\boldsymbol{a}+\mathrm{i}\,(\boldsymbol{c}-\boldsymbol{c}^{\dagger})=0$. The resulting solution is $u^{\prime}=-e^{-\mathrm{i}\,(k^{-1}x-k\,t)}\,\sigma_{2}\,a_{2}^{\dagger}\,\Omega^{-1}a_{1}$. $\Box$
Example 4.8.
Let $n=2$ and $\Gamma=\mathrm{diag}(\gamma,-\gamma^{\ast})$ with $\mathrm{Re}(\gamma)\neq 0$. Then we have
$\eta_{1}=\left(\begin{array}[]{cc}e^{-\varphi(\gamma)}&0\\ 0&e^{\varphi(\gamma^{\ast})}\end{array}\right)\,a_{1}\,,\qquad\tilde{\eta}_{2}=\left(\begin{array}[]{cc}e^{\varphi(\gamma)}&0\\ 0&e^{-\varphi(\gamma^{\ast})}\end{array}\right)\,a_{2}\,.$
Hermiticity of the matrices $\boldsymbol{a}$ and $\boldsymbol{b}$ means that the components $\boldsymbol{a}_{11},\boldsymbol{a}_{22},\boldsymbol{b}_{11},\boldsymbol{b}_{22}$ are real and $\boldsymbol{a}_{21}=\boldsymbol{a}_{12}^{\ast},\boldsymbol{b}_{21}=\boldsymbol{b}_{12}^{\ast}$. The off-diagonal parts of the Lyapunov equation amount to the constraint
$\boldsymbol{b}_{12}=\mathrm{i}\,\gamma\,\boldsymbol{a}_{12}\,$,
and the diagonal parts determine the diagonal components of $\Omega$,
$\Omega_{11}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\big(\mathrm{i}\,\gamma^{\ast}\boldsymbol{a}_{11}\,e^{-2\mathrm{Re}(\varphi(\gamma))}+\boldsymbol{b}_{11}\,e^{2\mathrm{Re}(\varphi(\gamma))}\big)\,$,
$\Omega_{22}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\big(\mathrm{i}\,\gamma\,\boldsymbol{a}_{22}\,e^{2\mathrm{Re}(\varphi(\gamma))}-\boldsymbol{b}_{22}\,e^{-2\mathrm{Re}(\varphi(\gamma))}\big)\,.$
The differential equations for $\Omega$ are now equivalent to
$\Omega_{12x}=-\frac{\mathrm{i}\,\boldsymbol{a}_{12}}{\gamma}\,,\quad\Omega_{21x}=\frac{\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}}{\gamma^{\ast}}\,,\quad\Omega_{12t}=\mathrm{i}\,\gamma\,\boldsymbol{a}_{12}\,,\quad\Omega_{21t}=-\mathrm{i}\,\gamma^{\ast}\boldsymbol{a}_{12}^{\ast}\,$,
which integrates to
$\Omega_{12}=-\mathrm{i}\,\boldsymbol{a}_{12}\,(\gamma^{-1}x-\gamma t)+c_{12}\,,\qquad\Omega_{21}=\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}\,(\gamma^{\ast-1}x-\gamma^{\ast}\,t)+c_{21}\,$,
with complex constants $c_{12}$ and $c_{21}$. The remaining condition (4.5) becomes
$c_{21}=c_{12}^{\ast}+\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}\,.$
Inserting the results in (4.8) yields the following solution of the $m_{2}\times m_{1}$ matrix FL equation,
$u^{\prime}=-\frac{1}{\mathrm{det}(\Omega)}\,\sigma_{2}\,a_{2}^{\dagger}\left(\begin{array}[]{cc}e^{-2\mathrm{i}\,\mathrm{Im}(\varphi(\gamma))}\,\Omega_{22}&-e^{2\varphi(\gamma^{\ast})}\,\Omega_{12}\\ -e^{-2\varphi(\gamma)}\,\Omega_{21}&e^{-2\mathrm{i}\mathrm{Im}(\varphi(\gamma))}\,\Omega_{11}\end{array}\right)a_{1}\,$,
where
$\mathrm{det}(\Omega)$ $=$ $\frac{1}{4\mathrm{Re}(\gamma)^{2}}\Big(\mathrm{i}\,\gamma\,\boldsymbol{a}_{22}\boldsymbol{b}_{11}\,e^{4\mathrm{Re}(\varphi(\gamma))}-\mathrm{i}\,\gamma^{\ast}\boldsymbol{a}_{11}\boldsymbol{b}_{22}\,e^{-4\mathrm{Re}(\varphi(\gamma))}-|\gamma|^{2}\boldsymbol{a}_{11}\boldsymbol{a}_{22}-\boldsymbol{b}_{11}\boldsymbol{b}_{22}\Big)$
$-\big|\mathrm{i}\,\boldsymbol{a}_{12}(\gamma^{-1}x-\gamma t)-c_{12}\big|^{2}-|\boldsymbol{a}_{12}|^{2}\rho-\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}c_{12}\,.$
Besides the exponential dependence on the independent variables via $e^{\varphi(\gamma)}$, there is also a rational dependence on $\gamma^{-1}x-\gamma t$ and its complex conjugate. Fig. 1 shows plots of the absolute values for two solutions of the scalar FL equation ($m_{1}=m_{2}=1$) from this class.
Figure 1: Plots of the absolute value of solutions of the scalar FL equation (with $\sigma_{1}=\sigma_{2}=1$) obtained with an anti-conjugate pair of spectral parameters from the family in Example 4.8. We set $\gamma=1+\mathrm{i}$, $a_{1}=(1,1)^{T}$, $a_{2}=(\mathrm{i}\,\gamma,1)^{T}$ and $c_{12}=1$ (left plot), respectively $c_{12}=50$ (right plot).
$\Box$
@@
where $\varphi$ stands for $\varphi(\gamma)$, and
$\rho=\frac{1}{2}\big(\gamma^{-2}x-t\big)\,$,
which introduces the same dependence on the independent variables that we met in the preceding example. After having determined
@@
If $\kappa:=2\mathrm{Re}(\gamma)\neq 0$, the solution of the Lyapunov equation is found from Example A.2 in Appendix Appendix A: Some facts about the Lyapunov equation to be
@@
Besides the exponential dependence via $A_{ij}$ and $B_{ij}$, the solution (4.16) also depends on the independent variables $x$ and $t$ through the linear expression $\rho$. Generically, the determinant of $\Omega$ has non-zero real and imaginary parts, so that $u^{\prime}$ is regular. We obtain matrix generalizations of “double pole solutions”, see [4] for the scalar case.
If $\gamma$ is imaginary, i.e., $\gamma=-\mathrm{i}\,k$ with $k\in\mathbb{R}\setminus\{0\}$, then we have $\rho=-\frac{1}{2}(x/k^{2}+t)=\rho^{\ast}$ and the Lyapunov equation requires (cf. Example A.4 in Appendix Appendix A: Some facts about the Lyapunov equation)
@@
$\mathrm{Im}(a_{21}\,a_{22}^{\ast})-\sigma_{1}\sigma_{2}\,\Big(\frac{1}{2}|a_{11}|^{2}-k\,\mathrm{Im}(a_{11}\,a_{12}^{\ast})\Big)=0\,.$ (4.20)
@@
so that $\mathrm{Re}(c_{22})$ and $\mathrm{Im}(c_{12})$ remain as free parameters. Let $k$ be determined by (4.19). The constants $a_{ij}$, $i,j=1,2$, are still constrained by (4.20). Then (4.16) is a quasi-rational solution of the $m_{2}\times m_{1}$ matrix FL equation. If $\boldsymbol{a}_{11}\neq 0$, for constant $t$ (respectively $x$) the numerators of components of $u^{\prime}$ grow at most with $x^{3}$ (respectively $t^{3}$), whereas $\det(\Omega)$ grows with the fourth power. On a line, along which $\rho$ and thus $x+k^{2}t$ is constant, $|u^{\prime}|\sim 1/|t|$ as $t\to\pm\infty$. As a consequence, the solution must be localized along a worldline, which is not a straight line. In particular, our results supplement the treatment of the scalar FL equation ($m_{1}=m_{2}=1$) in Section 4.1 of [4], also see Fig. 2 for two examples.
@@
The preceding examples exhaust the possibilities of solutions up to $n=2$. Mostly, they also present corresponding extensions to arbitrary $n$, but in the case, where $\Gamma$ is a Jordan block, we restricted our computations to $n=2$. The step to higher $n$ is straightforward, but the resulting expressions lengthy. According to Remark 4.4, we can superpose any number of solutions from the classes elaborated in the above examples.